OpenStudy (anonymous):

@greencat In one basketball season, Tim made 75% of the free throws he attempted. Tim made 36 free throws. How many free throws did he attempt? A. 27 B. 48 C. 61 D. 75

OpenStudy (anonymous):

I can help you get started \[36divx = 75\div100\]

OpenStudy (anonymous):

0.75

OpenStudy (kendricklamar2014):

27/36 = 75% 48/36 = 133% 61/36 = 169% 75/36 = 208%

OpenStudy (anonymous):

thanks

OpenStudy (kendricklamar2014):

Your Welcome

OpenStudy (anonymous):

@KendrickLamar2014 I thought 36 is HOW MANY free throws tim made, not the total he attempted. The question asks how many did he attempt. 36 was 75% of his total throws. Unless I'm reading wrong..

OpenStudy (kendricklamar2014):

Yeah your right ^

OpenStudy (kendricklamar2014):

In that case B.48 is the answer @SuperNinjaPower

OpenStudy (kendricklamar2014):

Thanks for catching my error @Everkids

OpenStudy (anonymous):

@KendrickLamar2014 Np lol, confused me big time for a minute there :3

OpenStudy (kendricklamar2014):

Yeah lol